How many Co atoms are there in a 11.8 gram sample of Co ?
1
Expert's answer
2018-03-06T07:41:33-0500
Number of moles of cobalt n = m/MW = 11.8 g/59 g/mole = 0.2 moles Co
Number of Co atoms then N = Na x n = 6.02 x 10^23 atoms/mole x 0.2 moles = 1.2 x 10^23 atoms